Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the company’s products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 3,200 helmets, using 1,984 kilograms of plastic. The plastic cost the company $15,078.

According to the standard cost card, each helmet should require 0.54 kilograms of plastic, at a cost of $8.00 per kilogram.

Required:
1.

According to the standards, what cost for plastic should have been incurred to make 3,200 helmets? How much greater or less is this than the cost that was incurred? (Round Standard kilograms of plastic per helmet to 2 decimal places.)

Number of helmets:

STandard kilograms of plastic per helmet:

Total standard kilograms:

standard cost per kilogram:

Total standard cost:

Actual cost incurred:

Total standard cost:

Total material variance:

     


2.

Break down the difference computed in (1) above into a materials price variance and a materials quantity variance. (Round your actual materials price to two decimal places, and round your final answers to the nearest whole dollar. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

Materials price variance:

    Materials quantity variance:
Answer #1

1)

Number of helmets: 3200
STandard kilograms of plastic per helmet: 0.54
Total standard kilograms: 1728
standard cost per kilogram: 8
Total standard cost: 13824
Actual cost incurred: 15078
Total standard cost: 13824
Total material variance 1254 U

Material price variance == (8*1984-15078) = 794 F

Material quantity variance = (1728-1984)*8 = 2048 U